OpenStudy (anonymous):

Please help What is the solution set of |2x – 3| – 8 = –1? 7 or –7 5 or –2 5 or 2 No solution

OpenStudy (nurali):

|2x – 3| – 8 = –1 |2x – 3| = 7 2x – 3=7 or 2x – 3=-7 2x =10 or 2x = -4 x=5 or x=-2